OpenStudy (anonymous):

help! (5)/(x+4) average rate of change from 1 to x. so far i got this equation [(5)/(x+4)]-1(1/(x-1)

OpenStudy (phi):

Is the question, if we have this function: \[ y= \frac{5}{x+4} \] what is its average rate of change between 1 and x ? If so, you want to find the change in y divided by the change in x when x= 1 y= 5/(1+4)= 5/5 =1 when x is x y = 5/(x+4) the change in y is \[ \frac{5}{x+4} - 1= \frac{5}{x+4}- \frac{(x+4)}{x+4} = \frac{5-x-4}{x+4}= \frac{1-x}{x+4}\] the change in x is x-1 the average rate is change in y divided by change in x can you finish ?

OpenStudy (anonymous):

is it x+4?

OpenStudy (anonymous):

or 1/x+4

OpenStudy (anonymous):

? /:

OpenStudy (amistre64):

if im reading it correctly, the avg rate of change is the slope between f(x) and f(1) \[\frac{f(x)-f(1)}{x-1}\]

OpenStudy (amistre64):

it looks like phi determined that as well

OpenStudy (anonymous):

yeah. so far i have (1-x)/(x+4) all over (x-1) but i don't know how to simplify. i think it's 1/x+4

OpenStudy (amistre64):

\[\frac{a/b}{c}=\frac{a}{bc}\]

OpenStudy (amistre64):

you can pull out a negative it seems to cancel like terms

OpenStudy (amistre64):

\[\frac{(1-x)}{(x-1)(x+4)}\] \[\frac{\cancel{(1-x)}}{-\cancel{(1-x)}(x+4)}\] \[\frac{1}{-(x+4)}\]

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

youre welcome